OpenStudy (ksaimouli):

find all solutions for 2cosx sinx-cosx=0

OpenStudy (amistre64):

we can factor out a cosx

OpenStudy (amistre64):

cos(x) (2sin(x)-1) = 0 when cos(x) = 0 when 2sin(x) = 1

OpenStudy (amistre64):

these trigs for "all solutions" tend to have to have a n*pi or some such in it to account for the periodicity

OpenStudy (ksaimouli):

0,2pie

OpenStudy (amistre64):

i take it that means we got an interval of 0 to 2pi :)

OpenStudy (ksaimouli):

ya

OpenStudy (amistre64):

cos(x) = 0, when we are at 90 and 270 degrees sin(x) = 1/2 when x= 60 and 120 if i remember it right

OpenStudy (amistre64):

sin is 30 and 150 .... those always blur together for me

OpenStudy (ksaimouli):

but the answ are pie/6,pie/2,5pie/6,3pie/2

OpenStudy (amistre64):

yeah, convert the degrees to radians and your set

OpenStudy (amistre64):



OpenStudy (ksaimouli):

can u plz give me the answer in radians

OpenStudy (ksaimouli):

y ai got u thanks